# Reverse Modeling in Large Language Models

### v1

@@ Abstract
[…]
utf8
@@ 1 Introduction
LLMs Touvron et al. (2023); Jiang et al. (2023) have shown impressive capabilities across diverse natural language processing tasks and beyond. These capabilities are primarily attributed to the learning of extensive corpora that cover general world knowledge Kaplan et al. (2020). These corpora are created in human society and often demonstrate human bias, including the inherently forward-oriented human cognition Bergen and Chan (2005); De Kerckhove and Lumsden (2013), e.g., reasons may precede outcomes and solutions can be deduced from given information in most cases of the grad school math dataset Mitra et al. (2024). In contrast, reverse thinking presents more cognitive challenges due to its contradiction with innate commonsense and human logic. It inspires us to explore the question of can LLMs do reverse modeling or will they face similar challenges as humans?, and could reverse modeling benefit the learning of LLMs?. To study this, we simulate reverse-modeling data by directly reversing entire paragraphs or documents at the token level. Please note that this is the simplest and extreme way but may not be the optimal way of emulating reverse thinking. We train LLMs with these simulated texts and conduct a comprehensive analysis. Overall results indicate that LLMs learn forward- and reverse-modeling texts equally well when trained from scratch. However, performance varies across text samples. Some are suited to reverse modeling, while others favor forward modeling. Notably, we found that texts suited for reverse modeling are often high-quality and more logically coherent. Training on them, the original “forward-modeling” LLMs can be improved. We conduct the empirical validation on language understanding benchmarks, such as Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020). In summary, this paper has two main contributions. First, it demonstrates and analyzes the performance of LLMs learned from forward- and reverse-modeling texts. Second, it shows that the data selected based on the losses of forward- and reverse-modeling texts can further improve LLMs.
@@ 2 Related Work
In this paper we utilize the reverse text for model training. Previous work on reverse inputs falls into three main areas. The first area involves using reverse text in machine translation. Studies show that using decoders to process text both left-to-right and right-to-left within an encoder-decoder framework improves machine translation performance Zhou et al. (2019); Gu et al. (2019), a finding later extended to LLMs Nguyen et al. (2024). Concurrently, Wu et al. (2018) examines the relationship between error propagation and reverse direction decoding in machine translation. The second area focuses on the reversal curse Berglund et al. (2023); Zhu et al. (2024), where an LLM trained to understand “A is B” may struggle to generalize to “B is A”. Reversing the text is proposed as a solution to this problem Golovneva et al. (2024); Guo et al. (2024). These two streams of works focus on machine translation or the reversal curse. Third, a recent work Papadopoulos et al. (2024) also explores the direction of input text, but there are two key differences compared to ours: (1) Our work is inspired by the concept of reverse thinking, while the reversed input is one simulating solution; (2) We further analyze it across different domains and inference steps and discover a valuable tool for assessing data quality. Our applications are partially related to the training data selection for LLMs, which is mainly divided into heuristic and model-based methods Longpre et al. (2023). Heuristic methods filter out low-quality data by defining various rules, such as the ratio of nouns and verbs Raffel et al. (2020); Penedo et al. (2023); Chowdhery et al. (2023); Sharma et al. (2024). Model-based methods filter data by training selection models or based on the perplexity of language models Wenzek et al. (2019); Xie et al. (2023); Wettig et al. (2024). However, our data selection method is an extra bonus derived from the reverse modeling analysis.
@@ 3 Experimental Settings
In Research Question (RQ) 1, we used the multilingual mC4111English, German, Korean, Arabic from https://huggingface.co/datasets/allenai/c4 Raffel et al. (2020) dataset to compare LLMs’ ability to handle forward and reverse texts under continued and from-scratch pretraining settings. In subsequent experiments, we used the English SlimPajama [^2] We use the widely-used public sampled version for experiments: https://huggingface.co/datasets/DKYoon/SlimPajama-6B Soboleva et al. (2023a) dataset, which includes seven different source domains. Testing LLMs trained on the multilingual mC4 dataset with samples from the SlimPajama dataset can further confirm our findings are general. More details are in Appendix A.
@@ 4 Experiments
RQ1: Can LLMs do reverse modeling?
@@
To explore LLMs’ reverse modeling capabilities, we investigate two pre-training approaches: (1) continued training from a well-trained model checkpoint and (2) pretraining from scratch with random initialization. Specifically, we train models fed with forward input and reverse text using the two approaches separately. Figure 1 compares training losses (average sample losses within training batches) for English using both methods on the mC4 dataset, while Figure 5 in the Appendix shows analogous results for other languages. In the continued pretraining setting, the forward loss for forward-modeling remains stable due to extensive training in the initial pretraining stage. In contrast, the reverse loss for reverse modeling, initially high, decreases rapidly after a few training steps. Notably, the forward loss is consistently lower than the reverse loss during continued pretraining. We speculate this occurs because the initial pretraining corpora consists entirely of forward-direction texts, imparting a natural directional bias to the LLMs. Consequently, the models find processing reverse information more challenging, similar to human difficulties with reverse thinking. Interestingly, in the from-scratch pretraining, the loss curves for both text directions converge almost identically. This pattern, also observed in other languages, indicates that LLMs can learn to process forward and reverse-modeling inputs with similar proficiency. This is because the model learns from both forward and reverse texts simultaneously with randomly initialized parameters, avoiding the initial forward-direction bias in well-trained models. RQ2: Does data domain influence the ability of LLMs’ reverse modeling? Based on the observation in RQ1, we focus on the from-scratch pretraining setting, where trained LLMs show almost equal losses from both forward and reverse directions. This raises the question of whether reverse loss consistently equals forward loss across all texts or if there are instances where reverse learning incurs a lower or higher loss. To explore this, we use the Slimpajama Soboleva et al. (2023b) text dataset, which covers a broad range of domains, for case-level evaluation. We first compute the average loss difference (Forward Loss - Reverse Loss) for each text and associate each text with its corresponding data source label. The overall case-level loss difference distribution across different source domains is shown in Figure 2. Observed that the loss differences of the text samples are centered around zero, showing an approximately normal distribution. Importantly, this indicates that reverse-direction loss is not universally higher than forward-direction loss. In fact, for over half of the texts, reverse prediction of the next tokens is comparatively easier.
Figure 2: Loss difference distribution across domains.
As indicated in Figure 2, compared to web-scraped corpora such as Wikipedia and Common Crawl, the distributions of loss differences from Book and ArXiv are generally less skewed towards easier forward-modeling. Furthermore, a larger proportion of texts in Book and ArXiv are easier to predict in the reverse direction compared to the forward direction. Considering that texts from books and academic papers are typically of higher quality than web-scraped texts, we speculate that texts, where reverse prediction is more effective, are likely to exhibit better logical coherence and fluency, indicative of their superior quality. This conjecture is also reflected in domains related to code, StackExchange, and Github. From the perspective of natural language, code often features monotonous syntax and repetitive vocabulary. From the perspective of human forward thinking and its reflection in written texts, the forward-direction prediction task, which involves predicting the future from the present, is inherently more challenging. Conversely, the reverse-direction token prediction operates from known outcomes back to their origins, potentially simplifying the task. RQ3: What features make texts easier to process in the reverse direction?
Figure 3: Mean step loss of full data during decoding.
Figure 4: Mean step loss of selected texts with the Top-$10\%$ and Bottom-$10\%$ loss differences during decoding.
@@ 5 Conclusions
In conclusion, our results demonstrate that LLMs can learn from both forward and reverse-modeling texts with comparable proficiency when trained from scratch, where human always struggle with reverse texts. This study also highlights the potential benefits of incorporating training data that favors reverse modeling. Our findings underscore the importance of exploring diverse cognitive frameworks to enhance the capabilities of LLMs.
6 Limitations
While our study demonstrates promising results in training LLMs with reverse modeling, several limitations should be acknowledged to provide a comprehensive understanding of the findings and guide future research. Firstly, the simulation of reverse modeling by simply reversing token sequences may not fully capture the complexity and nuances of true reverse cognitive processes. This approach reduces reverse modeling to a syntactic level, potentially overlooking deeper semantic and contextual factors intrinsic to human reverse modeling. Secondly, the evaluation metrics used in our study, such as performance on downstream benchmarks like MMLU, may not fully encompass the benefits or limitations of reverse modeling training. These metrics primarily measure specific aspects of language understanding and reasoning, potentially overlooking other critical dimensions influenced by reverse modeling, such as creativity or problem-solving skills. Lastly, our research does not address the potential computational and resource challenges associated with training LLMs on reverse modeling texts. The increased complexity and processing demands could pose significant barriers to practical applications, particularly in resource-constrained environments. In conclusion, while our findings offer valuable insights into the potential of reverse modeling in LLMs, addressing these limitations is crucial for advancing this line of research. Future studies should aim to develop more sophisticated methods for simulating reverse modeling, explore diverse and naturally occurring datasets, and consider a broader range of evaluation metrics to fully understand and harness the benefits of reverse cognitive processes in LLMs.
@@ A Implementation Details
In our initial experiments, we explore the effects of a higher learning rate ($2e^{-4}$) and extend training epochs ($2$ epochs) under the continued training setting. This exploration is based on the assumption that reverse modeling might need more epochs or a higher learning rate than forward modeling to overcome the pre-trained directional bias. While this increases the convergence speed, the final loss is nearly the same as when using a lower learning rate or a single training epoch. Thus, to ensure consistency in our comparisons, we fix the learning rate as $5e^{-5}$ and set the batch size to $48$ (with each batch consisting of $48$ paragraphs). Following established practices in LLM training Touvron et al. (2023); Chowdhery et al. (2023), we train for one epoch in all experiments. The number of training steps depends on the dataset size and batch size. For instance, training on a dataset with $1$ billion tokens requires approximately $2,400$ steps with our hyperparameter settings. We use Llama2-7B Touvron et al. (2023) as the LLM backbone for Research Questions 1–3 (Section 4). All experiments are conducted using $8$ NVIDIA A100-SXM-80GB GPUs, and the application experiments in Section 4 also use the Llama2-7B model.

### v2

@@ Abstract
Humans are accustomed to reading and writing in a forward manner, and this natural bias extends to text understanding in auto-regressive large language models (LLMs). This paper investigates whether LLMs, like humans, struggle with reverse modeling, specifically with reversed text inputs. We found that publicly available pre-trained LLMs cannot understand such inputs. However, LLMs trained from scratch with both forward and reverse texts can understand them equally well during inference across multiple languages. Our case study shows that different-content texts result in different losses if input (to LLMs) in different directions—some get lower losses for forward while some for reverse. This leads us to a simple and nice solution for data selection based on the loss differences between forward and reverse directions. Using our selected data in continued pretraining can boost LLMs’ performance by a large margin across different language understanding benchmarks.
@@ 1 Introduction
LLMs Touvron et al. (2023); Jiang et al. (2023) have shown impressive capabilities in various natural language processing tasks and beyond. These capabilities are primarily attributed to the learning of extensive corpora that cover general world knowledge Kaplan et al. (2020). These corpora are created in human society and often demonstrate human bias, including inherently forward-oriented human cognition Bergen and Chan (2005); De Kerckhove and Lumsden (2013), e.g., reasons may precede outcomes and solutions can be deduced from given information in most cases of the grad school math dataset Mitra et al. (2024). In contrast, reverse thinking presents more cognitive challenges due to its contradiction with innate common sense and human logic Chen et al. (2024). It inspires us to explore the following questions:
  • Can LLMs perform reverse modeling or will they face similar challenges as humans?
  • Can reverse modeling benefit the learning of LLMs?
To study this, we simulate reverse-modeling data by directly reversing entire paragraphs or documents at the token level. Please note that this is the simplest and extreme way, but may not be the optimal way of emulating reverse thinking. We train LLMs with these simulated texts and conduct a comprehensive analysis. Overall results indicate that LLMs learn forward- and reverse-modeling texts equally well when trained from scratch. However, performance varies across text samples. Some are suited to reverse modeling, while others favor forward modeling. Notably, we find that the texts suited for reverse modeling are often of high quality and more logically coherent. Training on them, the original “forward-modeling” LLMs can be improved. We perform empirical validation on language understanding benchmarks, such as Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020). In summary, this paper has two main contributions.
  • We examine how LLMs process and learn from text in both forward and reverse directions, demonstrating consistent patterns across multiple languages.
  • We show that strategically selecting training data based on the losses of forward- and reverse-modeling leads to improved model capabilities.
@@ 2 Related Work
In this paper we utilize the reverse text for model training. Previous work on reverse inputs falls into three main areas. The first area involves the use of reverse text in machine translation. Studies show that using decoders to process text both left-to-right and right-to-left within an encoder-decoder framework improves machine translation performance Zhou et al. (2019); Gu et al. (2019), a finding later extended to LLMs Nguyen et al. (2024). Concurrently, Wu et al. (2018) examines the relationship between error propagation and reverse direction decoding in machine translation. The second area focuses on the reversal curse Berglund et al. (2023); Zhu et al. (2024), where an LLM trained to understand “A is B” may struggle to generalize to “B is A”. Reversing the text is proposed as a solution to this problem Golovneva et al. (2024); Guo et al. (2024). These two streams of work focus on machine translation or the reversal curse. Third, a recent work Papadopoulos et al. (2024) also explores the direction of input text, but there are two key differences compared to ours: (1) Our work is inspired by the concept of reverse thinking, while the reversed input is one simulating solution; (2) We further analyze it across different domains and inference steps and discover a valuable tool for assessing data quality.
Our applications are partially related to the selection of training data for LLMs, which is divided mainly into heuristic and model-based methods Longpre et al. (2024). Heuristic methods filter out low-quality data by defining various rules, such as the ratio of nouns and verbs Raffel et al. (2020); Penedo et al. (2023); Chowdhery et al. (2023); Sharma et al. (2024). Model-based methods filter data by training selection models or based on the perplexity of language models Wenzek et al. (2020); Xie et al. (2023); Wettig et al. (2024). However, our data selection method is an extra bonus derived from the reverse modeling analysis.
@@ 3 Experimental Settings
In Research Question (RQ) 1, we used the multilingual mC4111English, German, Korean, Arabic from https://huggingface.co/datasets/allenai/c4 Raffel et al. (2020) dataset to compare LLMs’ ability to handle forward and reverse texts under continued and from-scratch pretraining settings. In subsequent experiments, we used the English SlimPajama [^2] We use the widely-used public sampled version for experiments: https://huggingface.co/datasets/DKYoon/SlimPajama-6B Soboleva et al. (2023) dataset, which includes seven different source domains. Testing LLMs trained on the multilingual mC4 dataset with samples from the SlimPajama dataset can further confirm our findings are general. More details are in Appendix A.
@@ 4 Experiments
RQ1: Can LLMs perform reverse modeling?
@@
To explore LLMs’ reverse modeling capabilities, we investigate two pre-training approaches: (1) continued training from a well-trained model checkpoint and (2) pretraining from scratch with random initialization. Specifically, we train models fed with forward input and reverse text using the two approaches separately. Figure 1 compares training losses (average sample losses within training batches) for English using both methods on the mC4 dataset, while Figure 7 in the Appendix B shows analogous results for other languages.
In the continued pretraining setting, the forward loss for forward-modeling remains stable due to extensive training in the initial pretraining stage. In contrast, the reverse loss for reverse modeling, initially high, decreases rapidly after a few training steps. Notably, the forward loss is consistently lower than the reverse loss during continued pretraining. We speculate this occurs because the initial pretraining corpora consists entirely of forward-direction texts, imparting a natural directional bias to the LLMs. Consequently, the models find processing reverse information more challenging, similar to human difficulties with reverse thinking.
Interestingly, in the from-scratch pretraining, the loss curves for both text directions converge almost identically. This pattern, also observed in other languages, indicates that LLMs can learn to process forward and reverse-modeling inputs with similar proficiency. This is because the model learns from both forward and reverse texts simultaneously with randomly initialized parameters, avoiding the initial forward-direction bias in well-trained models.
Figure 2: Loss difference distribution across domains.
RQ2: Does data domain influence the ability of LLMs’ reverse modeling?
Based on the observation in RQ1, we focus on the from-scratch pretraining setting, where trained LLMs show almost equal losses from both forward and reverse directions. This raises the question of whether reverse loss consistently equals forward loss across all texts or if there are instances where reverse learning incurs a lower or higher loss. To explore this, we use the SlimPajama Soboleva et al. (2023) text dataset, which covers a broad range of domains, for case-level evaluation.
Given a text sequence represented by tokens $\{V_{1},V_{2},\cdots,V_{N}\}$, for each position $t$ in the sequence ($0\leq t\leq N$), a LLM can generate a probability distribution over possible next tokens. We compute the cross-entropy loss at each position $t$, resulting in two sequences of loss values: $\{F_{1},F_{2},\cdots,F_{N}\}$ for the forward sequence and $\{R_{1},R_{2},\cdots,R_{N}\}$ for the reverse sequence.
We first compute the average loss difference (computed as $\frac{1}{N}(\sum_{i=1}^{N}F_{i}-\sum_{i=1}^{N}N_{i})$) for each text and associate each text with its corresponding data source label. The overall case-level loss difference distribution across different source domains is shown in Figure 2. Observed that the loss differences of the text samples are centered around zero, showing an approximately normal distribution. Importantly, this indicates that reverse-direction loss is not universally higher than forward-direction loss. In fact, for over half of the texts, reverse prediction of the next tokens is comparatively easier.
As indicated in Figure 2, compared to web-scraped corpora such as Wikipedia and Common Crawl, the distributions of loss differences from Book and ArXiv are generally less skewed towards easier forward-modeling. Furthermore, a larger proportion of texts in Book and ArXiv are easier to predict in the reverse direction compared to the forward direction. Considering that texts from books and academic papers are typically of higher quality than web-scraped texts, we speculate that texts, where reverse prediction is more effective, are generally more coherent, naturally flowing. Table 1 summarizes the randomly selected examples from the reverse easier and forward easier texts. The reverse easier texts display a coherent structure and smooth flow, making them easy for readers to follow. In contrast, the forward easier texts are relatively low-quality, less coherent, and often repetitive. This conjecture is also reflected in domains related to code, StackExchange, and Github. From the perspective of natural language, code often features monotonous syntax and repetitive vocabulary.
From the perspective of human forward thinking and its reflection in written texts, the forward-direction prediction task, which involves predicting the future from the present, is inherently more challenging. Conversely, the reverse-direction token prediction operates from known outcomes back to their origins, potentially simplifying the task.
RQ3: What features make texts easier to process in the reverse direction?
Figure 3: Assumptions on the step-by-step loss dynamics of full text data during decoding.
Figure 4: Empirical step-by-step loss dynamics of full text data during decoding.
To further validate our hypothesis, we conduct a detailed analysis of step-by-step loss changes during token decoding. While the aggregated view in RQ2 is informative, it hides the underlying step-by-step dynamics of the loss. Given $m$ text sequences in the SlimPajama dataset, we can obtain the two step-by-step loss sequences $\{\frac{\sum_{j=1}^{m}F_{1}^{j}}{m},\frac{\sum_{j=1}^{m}F_{2}^{j}}{m},\cdots,\frac{\sum_{j=1}^{m}F_{N}^{j}}{m}\}$ for the forward modeling and $\{\frac{\sum_{j=1}^{m}R_{1}^{j}}{m},\frac{\sum_{j=1}^{m}R_{2}^{j}}{m},\cdots,\frac{\sum_{j=1}^{m}R_{N}^{j}}{m}\}$ for the reverse modeling. We exclude the first and last tokens with step loss = 00 to avoid sharp changes at the start and end. To account for different text lengths, we normalize the steps of all texts to the interval $(0,1)$.
Given our findings that LLMs can effectively learn both forward and reverse modeling when trained from scratch, we initially hypothesize a straightforward relationship between the step-by-step loss of two directions, which is shown as assumption (a) in Figure 3. The assumption (a) is that forward and reverse modeling would exhibit similar loss patterns throughout the sequence, explaining the near-zero mean difference in average losses in RQ2. However, our empirical results, presented in Figure 4, reveal a more nuanced dynamic. The results instead support assumption (b) in Figure 3: reverse prediction becomes progressively more accurate as contextual information accumulates, while forward prediction maintains consistent difficulty levels across the sequence. These trajectories intersect at a critical intersection point, before which reverse prediction shows higher loss values, and after which it demonstrates lower loss values compared to forward prediction. Note that this pattern emerges consistently across all the texts. It is a statistical characteristic of all the texts in our datasets and is independent of text quality, representing a fundamental difference of LLM’s forward- and reverse-modeling behaviors.
To further understand this dyanmic, we analyze extreme cases (those in the top and bottom $10\%$ of average loss differences) to identify the features that drive these divergent patterns and to examine how these dynamics change in extreme cases. A straightforward hypothesis (assumption (c) in Figure 5) would suggest that extreme cases simply shift the reverse loss curve vertically while maintaining its shape, with top-$10\%$ cases shifting upward and bottom-$10\%$ cases shifting downward. Under this hypothesis, the intersection point between forward and reverse loss curves would show small distance changes. However, our findings in Figure 6 contradict this hypothesis and instead support assumption (d) in Figure 5: extreme cases primarily result in large horizontal shifts of the reverse loss dynamic, while the forward loss dynamic remains stable (simple vertical shift). In cases where forward loss substantially exceeds reverse loss (top-$10\%$ cases), we observe that reverse loss decreases rapidly, with the intersection point occurring very early in the sequence. Conversely, in cases where reverse loss is larger (bottom-$10\%$ cases), the intersection point is delayed until near the sequence end, with reverse loss consistently exceeding forward loss throughout most of the process.
Figure 5: Assumptions on the step-by-step loss dynamics of selected texts with the Top-$10\%$ and Bottom-$10\%$ loss differences during decoding.
Figure 6: Empirical step-by-step loss dynamics of selected texts with the Top-$10\%$ and Bottom-$10\%$ loss differences during decoding.
@@
The results shows that while the average loss difference is an aggregated metric, it effectively indicates different patterns in step-by-step loss dynamics. With our case studies in Table 1, we find the obvious text quality differences between the reverse-favoring cases and forward-favoring cases. This finding suggests that text quality is the key feature influencing the loss dynamics and the positions of intersection points. Our analysis also reveals a key insight: texts exhibiting early intersection points in their loss dynamics typically have higher loss differences and correspond to higher quality content. This relationship enables us to use the loss difference as a quality score for text quality assessment.
Application: Texts favoring reverse modeling can improve original LLMs.
@@
To further verify this assumption, we conduct continued pre-training on the publicly released Llama2-7B. Using the SlimPajama-6B Soboleva et al. (2023) as training data, we select 1B tokens with the lowest and highest $\mathcal{S}$ scores, respectively. The model’s performance is evaluated on MMLU Hendrycks et al. (2020). We also compare this with the following data selection strategies: (1) Random 1B: randomly sample 1B tokens, (2) Perplexity Lowest Ranked 1B: select the 1B tokens with the lowest perplexity by Llama2-7B.
The results from Table 2 show that the quality of training data significantly affects the performance of LLMs. Our high-quality data selection strategy ($\mathcal{S}$ Highest Ranked) outperforms other baselines, achieving the highest accuracy across various tasks on MMLU. Since the overall text quality of the SlimPajama 6B dataset is inferior to the text quality used in the pretraining of Llama2-7B, using the full 6B dataset does not improve over the original Llama2-7B. This suggests that the presence of low-quality data in unfiltered training sets degrades performance, as evidenced by the significant performance decline with low-quality selection strategy ($\mathcal{S}$ Lowest Ranked). This experiment supports the hypothesis that texts more effectively modeled by reversing are of higher quality and more beneficial for LLMs in acquiring world knowledge.
@@ 5 Conclusions
In conclusion, our results demonstrate that LLMs can learn from both forward and reverse-modeling texts with comparable proficiency when trained from scratch. This study also highlights the potential benefits of incorporating training data that favors reverse modeling. Our findings underscore the importance of exploring diverse reverse modeling frameworks to enhance the capabilities of LLMs.
Limitations
While our study demonstrates promising results in training LLMs with reverse modeling, several limitations should be acknowledged to provide a comprehensive understanding of the findings and guide future research.
Firstly, the simulation of reverse modeling by simply reversing token sequences may not fully capture the complexity and nuances of true reverse thinking processes. This approach reduces reverse modeling to a syntactic level, potentially overlooking deeper semantic and contextual factors intrinsic to human reverse modeling.
Secondly, the evaluation metrics used in our study, such as performance on downstream benchmarks like MMLU, may not fully encompass the benefits or limitations of reverse modeling. These metrics primarily measure specific aspects of language understanding and reasoning, potentially overlooking other critical dimensions influenced by reverse modeling, such as creativity or problem-solving skills.
Lastly, our research does not address the potential computational and resource challenges associated with training LLMs on reverse texts. The increased complexity and processing demands could pose significant barriers to practical applications, particularly in resource-constrained environments.
In conclusion, while our findings offer valuable insights into the potential of reverse modeling in LLMs, addressing these limitations is crucial for advancing this line of research. Future studies could aim to develop more sophisticated methods for simulating reverse modeling, explore diverse and naturally occurring datasets, and consider a broader range of evaluation metrics to fully understand and harness the benefits of reverse modeling in LLMs.
Acknowledgments
We would like to express our sincere gratitude to the anonymous reviewers for their insightful comments and valuable suggestions, which substantially improved this manuscript. S. Y. and J. W. are indebted to Xiaoqun Xiao, whose expertise in Arabic linguistics and textual analysis is instrumental in developing the multilingual aspects of this work.
@@ A Implementation Details
In our initial experiments, we explore the effects of a higher learning rate ($2e^{-4}$) and extend the training epochs ($2$ epochs) in a continued training setting. This exploration is based on the assumption that reverse modeling might need more epochs or a higher learning rate than forward modeling to overcome the pre-trained directional bias. While this increases the convergence speed, the final loss is nearly the same as when using a lower learning rate or a single training epoch.
Thus, to ensure consistency in our comparisons, we fix the learning rate as $5e^{-5}$ and set the batch size to $48$ (with each batch consisting of $48$ paragraphs). Following established practices in LLM training Touvron et al. (2023); Chowdhery et al. (2023), we train for one epoch in all experiments. The number of training steps depends on the size of the data set and the batch size. For instance, training on a dataset with $1$ billion tokens requires approximately $2,400$ steps with our hyperparameter settings.
We use Llama2-7B Touvron et al. (2023) as the LLM backbone for Research Questions 1–3 (Section 4). All experiments are conducted using $8$ NVIDIA A100-SXM-80GB GPUs, and the application experiments in Section 4 also use the Llama2-7B model.